Antioch Extraction, which mines ore in Montana, uses a calendar year for both financial-reporting and tax purposes. The following selected costs were incurred in December, the low point of activity, when 1,250 tons of ore were extracted:

Straight-line depreciation $ 27,500
Charitable contributions* 7,500
Mining labor/fringe benefits 200,000
Royalties 151,250
Trucking and hauling 201,195

*Incurred only in December.

Peak activity of 2,550 tons occurred in June, resulting in mining labor/fringe benefit costs of $408,000, royalties of $274,750, and trucking and hauling outlays of $256,195. The trucking and hauling outlays exhibit the following behavior:

Less than 1,250 tons $ 173,695
From 1,250–1,749 tons 201,195
From 1,750–2,249 tons 228,695
From 2,250–2,749 tons 256,195


Antioch uses the high-low method to analyze costs.

Required:
1. Classify the five costs listed in terms of their behavior: variable, step-variable, committed fixed, discretionary fixed, step-fixed, or semivariable.
2. Calculate the total cost for next February when 1,550 tons are expected to be extracted.
3-a. Is hauling 1,250 tons with respect to Antioch’s trucking/hauling cost behavior cost-effective?
3-b. If the company plans to extract 1,250 tons, at what number of tons can cost-effectiveness be achieved?
4. Distinguish between committed and discretionary fixed costs. If Antioch were to experience severe economic difficulties, which of the two types of fixed costs should management try to cut?
5. Speculate as to why the company’s charitable contribution cost arises only in December.

Ques )

Calculate the total cost for next February when 1,550 tons are expected to be extracted

Total cost      

If the company plans to extract 1,250 tons, at what number of tons can cost-effectiveness be achieved?

Number of tons      
Answer #1

Ans: As per HOMEWORKLIB RULESs, I can answer only four parts of the question.

1.

Straight-line depreciation $27500 Committed fixed
Charitable Contributions $7500 Discretionary fixed
Mining labor/Fringe benefits $200000 Variable
Royalties $151250 Semi-Variable
Trucking and Hauling $201195 Step fixed

Notes:

(a) Straight line depreciation remains fixed no matter the level of production and thus it is committed fixed costs.

(b) Charitable contribution is discretionary fixed costs as this cost is based on the decision of the user.

(c) As for the production between 1250-1749 tonnes, the trucking and haulage cost is $201195 and thus for 1250 tonnes it is 201195 and it is step-fixed cost. As the cost: is decided for every range.

(d) Calculations:

Mining labor/Fringe Benefits:

= $200000/1250 = $160/ton

= $408000/2550 = $160/ton

the cost is constant at the given volume levels that means it is a variable cost.

Royalties:

Variable Cost = ($274750 - $151250) / (2550 tonnes– 1250 tonnes)=$95/ton

Fixed Cost = 274750-(2550*95) = $32500

2.

Depreciation $27500
Charitable contributions 0
Mining labor (1550*$160) $248000
Royalties
Variable (1550*95) $147250
Fixed $32500
Trucking and hauling $201195
Total costs $656445

3a. The hauling of 1250 tons is not very cost-effective.

3b. It would be better if the company has decided to haul 1249 tons in order to decrease their costs by $27,500 ($201195-$17395)

4. Committed fixed costs are the fixed costs that arise from an entity’s ownership or due to the use of facilities and organizational structure. Examples of committed fixed costs include depreciation, property taxes, etc.

Discretionary fixed costs are the fixed costs which are incurred according the decision of the user. Generally these costs are incurred for the some specific purpose. Examples of discretionary fixed costs include charitable contributions, research and development.

In case of an economic difficulty, it is advisable to reduce discretionary fixed costs as it is possible to change these costs in the short run